If 38.7 g of a nonelectrolyte is dissolved into 218 g of water. The freezing point of the solution is measured to be -5.53 °C. Calculate the molar mass of the solute.

1
Expert's answer
2022-04-01T06:25:04-0400

ΔT = Kf * m

m – molality, moles of solute / kg solvent

moles = 38.7/M, M – molar mass

mass of water in kg: 218 g = 0.218 kg

So, 38.7/(M*0.218) is molality

ΔT = Kf * m

ΔT = (Kf * 38.7)/(M*0.218)

5.53 = (1.86 * 38.7) / (M*0.218)

M = 59.7 (g/mol)